Assume that stock market returns do not resemble a single-index structure. An investment fund analyzes 132 stocks in order to construct a mean-variance efficient portfolio constrained by 132 investments. They will need to calculate ____________ covariances. Multiple Choice 100 4,950 132 8,646
Answer #1

Number of covariances needed to construct mean-variance efficient portfolio=n*(n-1)/2=132*(132-1)/2=8646